1/3 ÷ (-4/7)(-1/6) + (-1/18)
Answer by fractalier(2101)

1/3 ÷ (-4/7)(-1/6) + (-1/18) =
1/3 + 4/42 + (-1/18) We need a lowest common denominator...here it's 126...so
1/3 + 4/42 + (-1/18) = 42/126 + 12/126 - 7/126 =
47/126